OpenStudy (caitlinnr14):

HELP PLEASE...MEDAL FOR CORRECT ANSWER (-1, 3) and perpendicular to y = -one thirdx + 7. y = 3x - 12 y = 3x + 6 y = 3x - 6 y = 3x

OpenStudy (anonymous):

I think it's the first one

OpenStudy (anonymous):

7 doesn't go into 6 nor 3

OpenStudy (ivettef365):

to be perpendicular the slope needs to be opposite reciprocal, and you have slope of -1/3, so the opposite reciprocal = 3 so use formula y = mx + b, substitute the (x,y) with (-1,3) to find value of b

OpenStudy (ivettef365):

so you have y = 3x + b 3 = 3(-1) + b what is b ??

OpenStudy (ivettef365):

???

OpenStudy (caitlinnr14):

6?

OpenStudy (ivettef365):

ok, so you have the equation m= 3 and b = 6 so you have y = 3x + 6

OpenStudy (caitlinnr14):

ok thank you!

OpenStudy (ivettef365):

yw :)